A solid weighs 0.850 kgf in air and 0.575 kgf in water. Find the density of the solid in SI units. Take density of water as 1000 kgm^–3
1
Expert's answer
2020-12-04T12:19:08-0500

"mg-\rho_1\cdot g\cdot V=0.85\cdot9.8=8.33\to \rho gV-\rho_1\cdot g\cdot V=8.33"


"mg-\rho_2\cdot g\cdot V=0.575\cdot9.8=5.64\to \rho gV-\rho_1\cdot g\cdot V=5.64"


"\frac{}{}" "\frac{\rho -\rho_1}{\rho -\rho_1}=\frac{8.33}{5.64}\to\rho=\frac{8.33\rho_2-5.64\rho_1}{8.33-5.64}="


"=\frac{8.33\cdot1000-5.64\cdot1.29}{8.33-5.64}=3094(kg/m^3)"